# Data-informativity conditions for structured linear systems with implications for dynamic networks

### v2

@@ Abstract
When estimating models of a multivariable dynamic system, a typical condition for consistency is to require the input signals to be persistently exciting, which is guaranteed if the input spectrum is positive definite for a sufficient number of frequencies. In this paper it is investigated how such a condition can be relaxed by exploiting prior structural information on the multivariable system, such as structural zero elements in the transfer matrix or entries that are a priori known and therefore not parametrized. It is shown that in particular situations the data-informativity condition can be decomposed into different MISO (multiple input single output) situations, leading to relaxed conditions for the MIMO (multiple input multiple output) model. When estimating a single module in a linear dynamic network, the data-informativity conditions can generically be formulated as path-based conditions on the graph of the network. The new relaxed conditions for data-informativity will then also lead to relaxed path-based conditions on the network graph. Additionally the new expressions are shown to be closely related to earlier derived conditions for (generic) single module identifiability.
@@
thanks: Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the European Research Council can be held responsible for them.thanks: This work was supported by VINNOVA Competence Center AdBIOPRO, contract [2016-05181] and by the Swedish Research Council through the research environment NewLEADS (New Directions in Learning Dynamical Systems), contract [2016-06079], and contract [2019-04956].thanks:   Version 20 November 2024.
@@ 1 Introduction
For consistently estimating a dynamic model of a linear multivariable system, a data-informativity condition needs to be satisfied, which in a prediction error framework is often formulated as a condition on the input spectrum matrix being positive definite at a sufficient number of frequencies (Söderström and Stoica, 1989). The condition is typically applied to situations where the inputs are general quasi-stationary signals with continuous spectral densities. For other types of inputs, as e.g. periodic signals, the concerned condition is conservative and replaced by conditions on the selection of a sufficient number of frequency components for each input and orthogonality of the signals over the different inputs (Pintelon and Schoukens, 2012). For particular black box model structures with finite order model sets, the spectrum condition can be further detailed, as shown e.g., in Gevers et al. (2008). The available theory for data informativity analysis seems to be mainly directed towards the situation of unstructured MIMO black box models, meaning that all entries in the MIMO model are parametrized. A relevant question that pops up is then: what are the consequences for the conditions for data-informativity if particular entries in the MIMO model are structurally zero, or non-zero but a priori known and therefore not necessarily parametrized?
@@
These structural situations can appear in the problem of identifying a single module in a dynamic network. For identifying a single module with the so-called direct method, it has been shown that it can be necessary to build a predictor model that has multiple inputs and multiple outputs (MIMO), (Ramaswamy and Van den Hof, 2021), in order to arrive at consistent module estimates. This is due to the necessity to deal with confounding variables, i.e., correlated disturbances between inputs and outputs of an estimation problem. The accompanying data-informativity conditions for estimating a single module in this MIMO predictor model, so far have been taken from a general non-structured MIMO estimation problem, i.e., requiring $\Phi_{w_{\!{\scriptscriptstyle\mathcal{D}}}}(\omega)\succ 0$ for a sufficient number of frequencies $\omega$, with $w_{\!{\scriptscriptstyle\mathcal{D}}}$ the set of inputs nodes in the predictor model. For a MISO setting, the data-informativity conditions have also been considered in Bombois et al. (2023) in the form of rank conditions on particular system-dependent matrices. For a generic satisfaction of the data-informativity conditions, very effective use has been made of a result of van der Woude (1991) and Hendrickx et al. (2019) where it is shown that rank conditions on transfer functions can generically be verified by path-based conditions on the graph of the network. This has led to the integration of path-based conditions in the selection of inputs and outputs of the predictor model to warrant that data-informativity conditions can be satisfied generically, (Van den Hof and Ramaswamy, 2020; Van den Hof et al., 2023).
In all these approaches, the data-informativity conditions for black box, fully parametrized models have been exploited. However, in the considered network problems structural information on the multivarible model might be readily available. The question that we address in this paper is therefore: can we exploit structural information on the system to be identified, in the form of a priori known entries in the transfer function matrix, for relaxing the data-informativity conditions? As a second and related topic, we investigate whether these conditions can be further relaxed in the case that we are actually only interested in consistently estimating a single module, rather than in a consistent full MIMO predictor model. Analysis of this second question will lead to two different sets of data-informativity conditions: one related to consistent single-module identification, and one related to consistent full predictor model identification. The importance of this distinction will also be discussed, and it will be illustrated that the path-based data-informativity conditions for consistent single module identification can be induced from the path-based single module identifiability results that were derived earlier Shi et al. (2022).
The remaining part of the paper is structured as follows. First in Sections 2 and 3, we will present the model setting and the notion of data-informativity in a prediction error framework. Then in Section 4 relaxed data-informativity conditions are presented for structured systems, and an example for illustrating the results is provided in Section 5. In Section 6 the conditions are formulated in terms of path-based conditions on the network graph. Consequences for the single module identification problem in dynamic networks are shown in Section 7. The paper concludes with another small example and some Conclusions.
@@ 2 The modeling setup
We consider the multivariable identification problem, where we consider a data generating dynamic system described by
@@
where output $y(t)\in\mathbb{R}^{p}$, input $u(t)\in\mathbb{R}^{m}$, the rational transfer function $G^{0}(\cdot)\in\mathbb{R}^{p\times m}(\cdot)$, $q$ the shift operator $qu(t)=u(t+1)$, $e(t)\in\mathbb{R}^{p}$ a multivariate white-noise process, and $H^{0}(\cdot)\in\mathbb{R}^{p\times p}(\cdot)$ a rational noise model, being monic, stable and stably invertible. For simplicity we will assume that $G^{0}(q)$ is stable too.
The data-generating system is modeled through a parametrized model $(G(q,\theta),H(q,\theta))$, with parameters $\theta\in\Theta\subset\mathbb{R}^{n}$. The model’s one-step ahead predictor can be defined as $\hat{y}(t|t-1;\theta):=\bar{\mathbb{E}}[y(t)|y^{t-1},u^{t}]$ where $u^{t}$ refers to the past of signal $u$ up to time instant $t$, and $\bar{\mathbb{E}}$ is the generalized expectation operator $\bar{\mathbb{E}}:=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{t=1}^{N}$ for quasi-stationary signals (Ljung (1999)).
@@
$\hat{y}(t|t-1;\theta)=(I-H(q,\theta)^{-1})y(t)+H(q,\theta)^{-1}G(q,\theta)u(t)$ (2)
which we write as
@@
In line with the corresponding definitions in the prediction error literature (Ljung (1999), Definition 8.1), we can now define the notion of data-informativity for the related predictor model.
@@
Consider a set of signals contained in $z$ and a predictor model
$\hat{y}(t|t-1;\theta)=W(q,\theta)z(t)$
for a parametrized set of models
$\mathcal{M}:=(G(q,\theta),H(q,\theta))_{\theta\in\Theta}.$
Then a quasi-stationary data set $Z^{\infty}:=\{z(t)\}_{t=0,\cdots,\infty}$ is informative enough with respect to the model set $\mathcal{M}$ if, for any two predictor models $W(q,\theta_{1})$ and $W(q,\theta_{2})$ in the model set, the equation
@@
implies that $W(e^{i\omega},\theta_{1})\equiv W_{2}(e^{i\omega},\theta_{2})$ for almost all $\omega$. $\Box$
Data-informativity guarantees that one can distinguish between different models in a model set, and consequently, that one can identify unique models on the basis of an experimental data set of infinite length. It is a necessary condition for estimating consistent models. If we do not constrain the order of the models in the model set $\mathcal{M}$, but allow $\mathcal{M}$ to consist of all linear time-invariant models of any order, then a typical sufficient condition for data-informativity is the condition that
$\Phi_{z}(\omega)\succ 0\ \ \ \mbox{for almost all }\omega.$ (5)
For open-loop situations where the signals $u$ and $e$ are uncorrelated, this can be further relaxed to the condition
@@
being referred to as the property of input signal $u$ being persistently exciting (Ljung (1999)).
@@ 3 Data-informativity conditions for consistent estimation
$\varepsilon(t,\theta)$ ::=:= $y(t)-\hat{y}(t|t-1;\theta),$ (8)
@@
and $Q$ any positive definite weighting matrix. Without loss of generality we will use $Q=I$ in the sequel.
@@
$\varepsilon(t,\theta)$ $=$ $H(q,\theta)^{-1}[G_{0}(q)-G(q,\theta)]u(t)+$ (9)
@@
$\varepsilon(t,\theta)=H(q,\theta)^{-1}[\Delta G(q,\theta)u(t)+\Delta H(q,\theta)e(t)]+e(t),$ (10)
with $\Delta G(q,\theta)=G^{0}(q)-G(q,\theta)$ and $\Delta H(q,\theta)=H^{0}(q)-H(q,\theta)$. The proof of consistency (Ljung, 1978) comes down to showing that
@@
    $(G(q,\theta),H(q,\theta))=(G^{0}(q),H^{0}(q))$ achieves a minimum of the cost function in (7), and
@@
    this minimum is unique.
If the transfer function $\Delta H(q,\theta)$ is strictly proper, which is guaranteed by the assumption that $H^{0}(q)$ and $H(q,\theta)$ both are monic, then it is straightforward to show that with $\Delta M(q,\theta):=\begin{bmatrix}\Delta G(q,\theta)&\Delta H(q,\theta)\end{bmatrix}$ and $\kappa(t):=\begin{bmatrix}u^{T}(t)&e^{T}(t)\end{bmatrix}^{T}$:
@@
With this expression it is obvious that condition (a) above is satisfied, reaching the minimum value of 00 of the cost function in (11) for $\Delta M=0$. For the uniqueness condition (b), we need to formulate a condition under which the following implication holds:
@@
[…]
@@ 4.1 Data-informativity for a single row / output
[…]
@@
While the vector $\kappa^{[j]}$ is composed of the signals related (the columns of) to nonzero terms in the corresponding model $[G(q,\theta)\ H(q,\theta)]_{j*}$, the theory actually holds for signals related to (the columns of) parametrized terms in the corresponding model. In other words, terms in the model that are known a priori, can be excluded from $\kappa^{[j]}$, as they do not serve as inputs to terms that are parametrized and need to be identified.

### v3

@@ Abstract
When estimating a single subsystem (module) in a linear dynamic network with a prediction error method, a data-informativity condition needs to be satisfied for arriving at a consistent module estimate. This concerns a condition on input signals in the constructed, possibly MIMO (multiple input multiple output) predictor model being persistently exciting, which is typically guaranteed if the input spectrum is positive definite for a sufficient number of frequencies. Generically, the condition can be formulated as a path-based condition on the graph of the network model. The current condition has two elements of possible conservatism: (a) rather than focussing on the full MIMO model, one would like to be able to focus on consistently estimating the target module only, and (b) structural information, such as structural zero elements in the interconnection structure or known subsystems, should be taken into account. In this paper relaxed conditions for data-informativity are derived addressing these two issues, leading to relaxed path-based conditions on the network graph. This leads to experimental conditions that are less strict, i.e. require a smaller number of external excitation signals. Additionally, the new expressions for data-informativity in identification are shown to be closely related to earlier derived conditions for (generic) single module identifiability.
@@
thanks: Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the European Research Council can be held responsible for them.thanks: This work was supported by VINNOVA Competence Center AdBIOPRO, contract [2016-05181] and by the Swedish Research Council through the research environment NewLEADS (New Directions in Learning Dynamical Systems), contract [2016-06079], and contract [2019-04956].thanks: Original version 20 November 2024; revised 29 July 2025.
@@ 1 Introduction
While the dynamic systems that are currently addressed in engineering and science, become more and more complex, large-scale and interconnected, the attention for modelling, monitoring and control aspects of interconnected systems, like e.g. dynamic networks, is growing. For the resulting control problems this has led to the development of decentralized, distributed and multi-agent type of control systems. On the modelling side, the development of analysis tools for systems over graphs has to be mentioned, and we have seen the development of a field called data-driven modeling (or system identification) in dynamic networks. This paper is situated in this latter area.
Identification of/in dynamic networks has been developed in a framework where dynamic systems are interconnected in a particular interconnection structure (topology), including cycles, that may be known a priori. In such a setting several different problems have been considered. Following the early contributions on uniqueness of network representations Goncalves&Warnick:08 and topology estimation Materassi&Innocenti:10 ; Materassi&Salapaka:12 , a framework for network identification in a prediction error setting is introduced in VandenHof&etal_Autom:13 , while subspace methods for network identification have been developed in Haber&Verhaegen_TAC:14 ; Verhaegen&etal:22 .
The vast literature on network identification problems in a prediction error setting, ranges from the problem of identifying the full network for a given topology Yuan:11 ; Weerts&etal_Autom:18_reducedrank ; Fonken&etal_Autom:22 , and the analysis and synthesis of identifiability conditions Weerts&etal_Autom:18_identifiability ; Hendrickx&Gevers&Bazanella_TAC:19 ; vanWaarde&etal_TAC:20a ; Legat&Hendrickx_CDC:20 ; Cheng&etal_TAC:22 ; Shi&etal_TAC:23 ; Mapurunga&etal:24 , to the problem of identifying the dynamics of a single module/link/subsystem in the network VandenHof&etal_Autom:13 ; Dankers&etal_TAC:16 ; Linder&Enqvist_ijc:17 ; Gevers&etal_sysid:18 ; Everitt&Bottegal&Hjalmarsson_Autom:18 ; Materassi&Salapaka:20 ; Ramaswamy&VandenHof_TAC:21 , which includes the problem of determining where to measure and where to excite the network, and the related data-informativity questions Bombois&etal_Autom:23 .
In this paper we will focus on this latter single module identification problem, in a setting where some dynamic modules in the network may be known a priori, while a single target module will need to be estimated consistently. We focus on the so-called local direct method (Ramaswamy&VandenHof_TAC:21, ) that is known to be able to exploit both measured external excitation signals and unmeasured disturbance signals for estimation. It has been shown (Ramaswamy&VandenHof_TAC:21, ) that it can be necessary to build a predictor model that has multiple inputs and multiple outputs (MIMO), in order to arrive at consistent estimates of a single target module. This is due to the necessity to deal with confounding variables, i.e., correlated disturbances between inputs and outputs of an estimation problem, a situation that can typically occur in dynamic networks. The resulting MIMO predictor model then will need to satisfy data-informativity conditions in order to arrive at consistent module estimates. The objective of this paper is to formulate path-based conditions, dependent on the a priori known graph of the network only, that guarantee data-informativity, and therefore will provide the conditions for consistent module estimates. These path-based conditions will refer to the locations of external signals (measured excitation signals and unmeasured disturbance signals) that enter the network.
@@
For specifying the role of a data-informativity condition, we need to go back to their origin, developed for classical (open-loop type of) identification problems in a prediction error setting. It is most often formulated as a condition on the input spectrum matrix being positive definite at a sufficient number of frequencies (Soderstrom&Stoica:89, ). The condition is typically applied to situations where the inputs are general quasi-stationary signals with continuous spectral densities. For other types of inputs, as e.g. periodic signals, the concerned condition is conservative and replaced by conditions on the selection of a sufficient number of frequency components for each input and orthogonality of the signals over the different inputs (Pintelon&Schoukens:12, ). For particular black box model structures with finite order model sets, the spectrum condition can be further detailed, as shown e.g., in Gevers&Bazanella&Miskovic_CDC:08 . The available theory for data informativity analysis seems to be mainly directed towards the situation of unstructured MIMO black box models, meaning that all entries in the MIMO model are parametrized. A relevant question that pops up is then: what are the consequences for the conditions for data-informativity if particular entries in the MIMO model are structurally zero, or non-zero but a priori known and therefore not necessarily parametrized? This latter situation is typically present in a dynamic network situation.
@@
In the current literature on dynamic network identification, the data-informativity conditions so far have been taken from a general non-structured MIMO estimation problem, i.e., requiring $\Phi_{w_{\!{\scriptscriptstyle\mathcal{D}}}}(\omega)\succ 0$ for a sufficient number of frequencies $\omega$, with $w_{\!{\scriptscriptstyle\mathcal{D}}}$ the set of inputs nodes in the predictor model. For a MISO setting, the data-informativity conditions have also been considered in Bombois&etal_Autom:23 in the form of rank conditions on particular system-dependent matrices. For a generic satisfaction of the data-informativity conditions, very effective use has been made of a result of vanderWoude:91 and Hendrickx&Gevers&Bazanella_TAC:19 where it is shown that rank conditions on transfer functions can generically be verified by path-based conditions on the graph of the network. This has led to the integration of path-based conditions in the selection of inputs and outputs of the predictor model to warrant that data-informativity conditions can be satisfied generically (VandenHof&Ramaswamy_CDC:20, ; VandenHof&etal_IFAC:23, ).
In all these approaches, the data-informativity conditions for black box, fully parametrized MIMO models have been exploited. However, in the considered network problem, structural information on the multivarible predictor model might be readily available (e.g. its topology, indicating which nodes are connected with which other nodes), and rather than consistently estimating the full MIMO predictor model, we can focus on consistently estimating the single target module only.
The question that we address in this paper is therefore: can we exploit structural information (topology) of the network, in the form of a priori known entries in the predictor model, to relax the data-informativity conditions and to focus these conditions on estimating a single target module only. In comparison with the current literature, these are two innovative steps to be taken: including the structural information, and addressing a single target module only.
We will analyse the situation by first addressing the structural information in a standard (open-loop type of) estimation problem in Sections 3-5, including a presentation of the modeling setup and the notion of data-informativity. After presenting an example in Section 6, we will then formulate the new results in terms of path-based conditions on the network graph in Section 7. With these new results, we can address the single module identification problem in dynamic networks in Section 8, with some examples collected in Section 9. The proofs of the most important results are collected in an Appendix. We start by presenting a very brief example.
2 A motivating example
In order to give a flavor of the type of results that will be achieved in this paper, we present an example network provided in Figure 1, where linear time-invariant systems $G_{jk}$ are interconnected in a network structure, and the node variables $w_{\ell}$ are the output signals of the corresponding summation points. $v_{k}$ are non-measured stationary stochastic disturbance signals, $v_{1}$ and $v_{2}$ being mutually correlated, and $x_{i}$ are possible locations for adding external signals. Our objective is to identify the target module $G_{12}$ from data $w_{\ell}$, $\ell=1,\cdots 6$, and possibly present measured $x_{i}$. In the current theory of the local direct method, a consistent estimation of $G_{12}$, can be achieved if the four external signals $x_{1}$, $x_{2}$, $x_{3}$ and $x_{5}$ are all present, with $x_{1}$ and $x_{2}$ being measured. Our new results, derived in this paper, will show that the presence of signals $x_{3}$ and $x_{5}$ only, with $x_{5}$ being measured, will be sufficient, thereby considerably reducing the experimental burden. This example will be treated in full detail in Section 9.2.
@@ 3 The modeling setup
For properly addressing the data-informativity problem, we start by considering a multivariable (open-loop) data-generating dynamic system described by
@@
where output $y(t)\in\mathbb{R}^{p}$, input $u(t)\in\mathbb{R}^{m}$, the rational transfer function $G^{0}(\cdot)\in\mathbb{R}^{p\times m}(\cdot)$, $q$ the shift operator $qu(t)=u(t+1)$, $e(t)\in\mathbb{R}^{p}$ a multivariate white-noise process, and $H^{0}(\cdot)\in\mathbb{R}^{p\times p}(\cdot)$ a rational noise model, being monic, i.e. $\lim_{z\rightarrow\infty}H(z)=I$, stable and stably invertible. For simplicity we will assume that $G^{0}(q)$ is stable too.
The data-generating system is modeled through a parametrized model $(G(q,\theta),H(q,\theta))$, with parameters $\theta\in\Theta\subset\mathbb{R}^{n}$. The model’s one-step ahead predictor can be defined as $\hat{y}(t|t-1;\theta):=\bar{\mathbb{E}}[y(t)|y^{t-1},u^{t}]$ where $u^{t}$ refers to the past of signal $u$ up to time instant $t$, and $\bar{\mathbb{E}}$ is the generalized expectation operator $\bar{\mathbb{E}}:=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{t=1}^{N}\mathbb{E}$ for quasi-stationary signals (Ljung:99 ).
@@
In line with the corresponding definitions in the prediction error literature (Ljung:99 , Definition 8.1), we can now define the notion of data-informativity for the related predictor model.
@@
Consider a set of signals contained in $z$ and a predictor model $\hat{y}(t|t-1;\theta)=W(q,\theta)z(t)$, as in (2) for a parametrized set of models $\mathcal{M}:=(G(q,\theta),H(q,\theta))_{\theta\in\Theta}$. Then a quasi-stationary data set $Z^{\infty}:=\{z(t)\}_{t=0,\cdots,\infty}$ is informative enough with respect to the model set $\mathcal{M}$ if, for any two predictor models $W(q,\theta_{1})$ and $W(q,\theta_{2})$ in the model set, the equation
@@
implies that $W(e^{i\omega},\theta_{1})\equiv W_{2}(e^{i\omega},\theta_{2})$ for almost all $\omega$, where $\|\cdot\|$ is the standard Euclidean norm. $\Box$
Data-informativity guarantees that one can distinguish between different models in a model set, and consequently, that one can identify unique models on the basis of an experimental data set of infinite length. It is a necessary condition for the consistent estimation of models. If we do not constrain the order of the models in the model set $\mathcal{M}$, but allow $\mathcal{M}$ to consist of all linear time-invariant models of any order, then a typical sufficient condition for data-informativity is the condition that $\Phi_{z}(\omega)\succ 0$ for almost all $\omega$. For open-loop situations where the signals $u$ and $e$ are uncorrelated, this can be further relaxed to the condition
@@
being referred to as the property of input signal $u$ being persistently exciting (Ljung:99 ).
@@ 4 Data-informativity conditions for consistent estimation
$\varepsilon(t,\theta)$ $:=$ $y(t)-\hat{y}(t|t-1;\theta),$ (6)
@@
and $Q$ any positive definite weighting matrix. Without loss of generality we will use $Q=I$ in the sequel, while for the consistency analysis we will assume that $(G^{0},H^{0})\in\mathcal{M}$.
@@
$\varepsilon(t,\theta)$ $=$ $H(q,\theta)^{-1}[G{\color[rgb]{0,0,0}{}^{0}}(q)-G(q,\theta)]u(t)+$ (7)
@@
$\varepsilon(t,\theta)=H(q,\theta)^{-1}\Delta M(q,\theta)\kappa(t)+e(t),$ (8)
with $\Delta M(q,\theta):=[\Delta G(q,\theta)\ \ \Delta H(q,\theta)]$, $\Delta G(q,\theta):=G^{0}(q)-G(q,\theta)$, $\Delta H(q,\theta):=H^{0}(q)-H(q,\theta)$ and $\kappa(t):=[u^{T}(t)\ \ e^{T}(t)]^{T}$. The proof of consistency (Ljung:78, ) comes down to showing that
@@
    $\Delta M(q,\theta)$ achieves the global minimum of the cost function in (5), and
@@
    this minimum is unique, i.e. it provides a single point estimate $(G(q,\theta^{*}),H(q,\theta^{*}))$.
If the transfer function $\Delta H(q,\theta)$ is strictly proper, which is guaranteed by the assumption that $H^{0}(q)$ and $H(q,\theta)$ both are monic, then it is straightforward to show that
@@
With this expression it is obvious that condition (a) above is satisfied, reaching the minimum value of 0 of the cost function in (9) for $\Delta M=0$. For the uniqueness condition (b), we need to formulate a condition under which the following implication holds:
@@
$\frac{1}{2\pi}\mbox{trace}\{\int_{-\pi}^{\pi}\Delta M(e^{i\omega},\theta)\Phi_{\kappa}(\omega)\Delta M^{T}(e^{-i\omega},\theta)d\omega\}=0$ (11)
@@ 5.1 Data-informativity for a single row / output
$\frac{1}{2\pi}\sum_{\ell=1}^{p}\{\int_{-\pi}^{\pi}\Delta M_{\ell}(e^{i\omega},\theta)\Phi_{\kappa}(\omega)\Delta M_{\ell}^{T}(e^{-i\omega},\theta)d\omega\}=0$ (12)
@@
While the vector $\kappa^{[j]}$ is composed of the signals related to (the columns of) nonzero terms in the corresponding model $G(q,\theta)\ H(q,\theta)]_{j*}$, the theory actually holds for signals related to (the columns of) parametrized terms in the corresponding model. In other words, terms in the model that are known a priori (like controllers in a network), can be excluded from $\kappa^{[j]}$, as they do not serve as inputs to terms that are parametrized and need to be identified.